# Unsupervised Post-Training for Multi-Modal LLM Reasoning via GRPO

### v1

@@ Abstract
Improving Multi-modal Large Language Models (MLLMs) in the post-training stage typically relies on supervised fine-tuning (SFT) or reinforcement learning (RL). However, these supervised methods require expensive and manually annotated multi-modal data–an ultimately unsustainable resource. While recent efforts have explored unsupervised post-training, their methods are complex and difficult to iterate. In this work, we are the first to investigate the use of GRPO, a stable and scalable online RL algorithm, for enabling continual self-improvement without any external supervision. We propose MM-UPT, a simple yet effective framework for unsupervised post-training of MLLMs. MM-UPT builds upon GRPO, replacing traditional reward signals with a self-rewarding mechanism based on majority voting over multiple sampled responses. Our experiments demonstrate that MM-UPT significantly improves the reasoning ability of Qwen2.5-VL-7B (e.g., 66.3$\rightarrow$72.9on MathVista, 62.9$\rightarrow$68.7on We-Math), using standard dataset without ground truth labels. MM-UPT also outperforms prior unsupervised baselines and even approaches the results of supervised GRPO. Furthermore, we show that incorporating synthetic questions, generated solely by MLLM itself, can boost performance as well, highlighting a promising approach for scalable self-improvement. Overall, MM-UPT offers a new paradigm for continual, autonomous enhancement of MLLMs in the absence of external supervision. Our code is available at https://github.com/waltonfuture/MM-UPT.
@@ 1 Introduction
Multi-modal Large Language Models (MLLMs) have achieved remarkable performance on a variety of vision-language tasks, ranging from image captioning to visual reasoning [21, 69, 19, 47, 42]. By combining the language understanding capabilities of large language models (LLMs) with visual perception, MLLMs can process and reason over both textual and visual information. The dominant paradigm for improving MLLMs in the post-training stage typically involves supervised fine-tuning (SFT) and reinforcement learning (RL). However, both SFT and RL rely on large volumes of high-quality and annotated multi-modal data, such as image captions, visual reasoning traces, ground truth answers, and human preference signals. As real-world tasks grow in complexity and quantity, a critical challenge emerges: curating and annotating high-quality data at scale becomes increasingly impractical. Thus, it is essential to explore new methods for improving MLLMs, such as using synthetic and unlabeled data.
@@
Previous works have studied the use of MLLMs themselves to generate synthetic instruction data for self-improvement through offline training techniques like SFT and DPO [64, 5, 55, 37, 38, 36]. These approaches typically involve complex pipelines with multiple stages, such as data generation, verification, and filtering, which are hard to iterate online. Fortunately, recent studies demonstrate notable success using online reinforcement learning (e.g. GRPO [34]) with verifiable rewards [10, 66, 28, 6] to enhance the reasoning capabilities of LLMs and MLLMs. A concurrent work, TTRL [71], further extends this line by applying GRPO on test-time scaling of LLMs. Thus, it is promising that online RL enables models to continuously improve, thus acquiring novel reasoning abilities that exceed corresponding base models’ capacity. Motivated by these insights, we aim to investigate a fundamental and open question: Can we enable MLLMs to continually and iteratively self-improve their reasoning abilities from fully unlabeled training data without any external supervision?
@@
To explore this question, we propose MM-UPT (Multi-Modal Unsupervised Post-Training), an easy-to-implement framework for unsupervised post-training in MLLMs. As shown in Figure 1, MM-UPT builds on GRPO [34], a stable and scalable online RL method that uses group-normalized rewards instead of explicit value functions. Unlike GRPO that relies on ground-truth labels to calculate rewards, our framework works by deriving implicit reward signals via majority voting over multiple sampled responses. In particular, majority voting aggregates multiple responses and selects the most frequent one, which has been widely used and shown effective to improve model performance [45, 40, 71]. Thus, we adopt the majority-voted answer to serve as a pseudo-label to calculate reward in our framework for unsupervised training, which encourages the model to prefer stable and high-consensus answers without any human annotated labels or external reward models.
@@
In our experiments, we focus on the domain of multi-modal reasoning, which is widely studied and inherently challenging. We explore two key scenarios for constructing unlabeled data assuming that labels are not available: (1) using human-created questions without ground-truth labels, and (2) employing synthetic questions generated by AI models, inherently lacking ground-truth labels. The first scenario is simulated by masking the answers in standard training datasets. For the second scenario, we employ two MLLM-driven strategies: (a) generating new questions from in-context examples (including the original image, question, and answer), and (b) generating questions based solely on the image. We evaluate MM-UPT across a range of reasoning benchmarks and observe significant performance improvements over the base models (e.g., 66.3$\rightarrow$72.9on MathVista, 62.9$\rightarrow$68.7on We-Math using Qwen2.5-VL-7B [1]). Our method also outperforms previous baseline methods, and is even competitive with supervised GRPO, underscoring the effectiveness of MM-UPT as a self-improving training strategy in a fully unsupervised manner. Additionally, we find that models trained on unlabeled synthetic data achieve performance competitive with those trained on the original unlabeled dataset, revealing strong potential for scalable self-improvement.
@@
  • We propose MM-UPT, a novel framework for unsupervised post-training of MLLMs that enables continuous self-improvement without requiring any external supervision.
@@
  • We investigate the use of synthetic data generated by the MLLM itself and find that training the MLLM on such data leads to notable performance gains. This reveals a promising path toward efficient and stable self-improvement in unsupervised post-training.
@@ 2 Related Works
Self Improvement. High-quality data obtained from human annotations has been shown to significantly boost the performance of LLMs across a wide range of tasks [16, 29, 10]. However, such high-quality annotated data may be exhausted in the future. This presents a substantial obstacle to the continual learning of advanced models. As a result, recent research has shifted toward self-improvement—leveraging data generated by the LLM itself without any external supervision [14, 61, 49, 26, 7, 38, 56, 71]. Several following works also explore unsupervised self-improvement in the multi-modal domain [9, 5, 55, 64, 36, 62]. Genixer [64] firstly introduces a comprehensive self-improvement pipeline including data generation and filtering for SFT. STIC [5] and SENA [37] construct preference data pairs for DPO [33] in a fully self-supervised manner. In contrast to these approaches which are complex and hard to iterate, the key distinction is that our method leverages online reinforcement learning using GRPO [10] at the post-training stage of MLLMs, which is more scalable, and supports continuous and iterative self-improvement without reliance on any external supervision. In addition, none of these previous methods focus on multi-modal reasoning tasks, which are considered challenging for current models.
Multi-modal Reasoning. Recently, the reasoning abilities of MLLMs have become a central focus of research [24, 70, 13]. In contrast to traditional LLM-based reasoning [25, 54, 10] that primarily relies on text, multi-modal approaches must both process and interpret visual inputs, significantly increasing the complexity of tasks such as geometric problem-solving and chart interpretation [2, 27, 63]. Several works in this field have sought to collect or synthesize a large scale of multi-modal reasoning data [58, 35, 30, 4]. Notably, the recent emergence of o1-like reasoning models [17] represents an initial step toward activating the slow-thinking capabilities of MLLMs, as demonstrated by several SFT-based methods, such as LLaVA-CoT [50], LLaVA-Reasoner [59], MAmmoTH-VL [11], and Mulberry [52]. Moreover, some concurrent works have further explored reinforcement learning approaches, particularly GRPO [34], in the post-training stage of MLLMs to enhance performance on multi-modal reasoning tasks [66, 28, 6, 44, 31]. While these supervised post-training methods have demonstrated promising results, our work explores a different direction by focusing on totally unsupervised post-training of MLLMs to self-improve the reasoning abilities.
@@ 3 The Framework of Multi-Modal Unsupervised Post-Training
Unlike traditional post-training techniques that require labeled data or external reward models, we propose MM-UPT (Multi-Modal Unsupervised Post-Training), a simple yet effective framework designed to operate purely on unlabeled training data. That is, the model must learn to self-improve without access to any external supervision such as ground-truth labels or additional reward models.
@@ 3.1 Problem Formulation
3.2 Method
To achieve this, MM-UPT introduces a self-rewarding mechanism using majority voting as pseudo-labels [45] based on the online reinforcement learning. In particular, MM-UPT is built upon the GRPO algorithm [34], which is widely used in the post-training stage of multi-modal LLMs. GRPO optimizes computational efficiency by eliminating the need for a separate value model; instead, it directly utilizes group-normalized rewards to estimate advantages. Specifically, for a question $q$ and the correlated image $I$ from the training dataset $Q$, GRPO samples a group of responses $O=\{o_{i}\}_{i=1}^{G}$ from the old policy $\pi_{old}$ and then optimizes the policy model by maximizing the following objective:
@@ 3.2 Method
Algorithm 1 MM-UPT: Multi-Modal Unsupervised Post-Training
@@
In this way, we compute pseudo-rewards via majority voting and apply standard GRPO to update the MLLM. This majority-based reward encourages the model to converge toward consistent, high-consensus responses, thereby enabling the model to further exploit its existing self-knowledge leveraging unlabeled data. The overview of our framework is shown in Figure 1 and Algorithm 1.
@@ 4 Experiments
Table 1: Main results on four multi-modal mathematical reasoning benchmarks. We report accuracy (%) for each method on MathVision, MathVerse, MathVista, and We-Math. All methods are conducted on the Qwen2.5-VL-7B backbone. MM-UPT outperforms other baseline methods, and is even competitive with supervised methods.
@@
+ MM-UPT Geometry3K 27.33 42.46 68.50 66.61 51.23
+ MM-UPT GeoQA 27.07 43.68 68.90 68.22 51.97
+ MM-UPT MMR1 26.15 44.87 72.90 68.74 53.17
@@ 4.1 Experimental Setup
Baseline Methods. Several prior works have explored self-improvement in both LLMs and MLLMs. Note that we focus on unsupervised self-improvement, we do not compare with methods that rely on external models (e.g., GPT-4o [16]) for supervision [60, 15, 67, 68, 55]. Instead, we compare with several totally unsupervised methods: LMSI [14], SRLM [49], Genixer [64], and STIC [5]. In particular, LMSI corresponds to supervised fine-tuning with self-generated content selected by majority voting. SRLM uses the model itself as LLM-as-a-Judge [65] to provide its own rewards during DPO [33] training. Genixer prompts the MLLM to first self-generate an answer and then self-check it. STIC applies DPO where original images and good prompts are used to generate preferred answers, and corrupted images and bad prompts to produce rejected answers. Additionally, we also compare with GRPO [34] and rejection sampling-based SFT [39], which are two strong supervised methods. The details of these baseline methods are shown in Appendix A.1.
@@
Implementation Details. We adopt the EasyR1 [53] framework for multi-modal unsupervised post-training, which is based on GRPO. Specifically, we set the training episodes to 15, and use AdamW optimizer [22] with a learning rate of $1\times 10^{-6}$, weight decay of $1\times 10^{-2}$, and gradient clipping at a maximum norm of 1.0. The KL divergence constraint $\beta$ in GRPO is set to 0.01 to stabilize the training. The vision tower of the multi-modal model is also tuned without freezing. Other hyperparameters follow the default settings provided in the EasyR1 framework.
4.2 Unsupervised Training on Standard Datasets
@@
Table 1 presents the main results on four challenging multi-modal mathematical reasoning benchmarks. We observe that MM-UPT achieves consistent improvements in average over the base Qwen2.5-VL-7B model across all datasets, also outperforming other baseline methods such as SRLM, LMSI, Genixer, and STIC. Notably, MM-UPT is able to improve the average score from 49.47 (base model) to 53.17 (with MMR1 dataset), demonstrating its effectiveness in leveraging unlabeled data for self-improvement. In comparison, previous baselines provide only marginal gains or even degrade performance on certain benchmarks, highlighting the limitations of existing methods when applied to already strong models in multi-modal reasoning tasks. Furthermore, we find that MM-UPT is even competitive with supervised post-training methods, such as rejection sampling-based SFT [39] and GRPO [34]. These results underscore the potential of MM-UPT to further exploit the knowledge embedded in multi-modal models for self-improvement.
4.3 Unsupervised Training on Synthetic Datasets
Table 2: Performance comparison of MM-UPT using different synthetic data generation strategies. Both “In-Context Synthesizing” and “Direct Synthesizing” approaches yield significant improvements over the base model and perform competitively with the “Original Questions” on average, demonstrating the effectiveness of synthetic data for unsupervised self-improvement.
@@ 4.3 Unsupervised Training on Synthetic Datasets
w/ Original Questions Geo3K 27.33 42.46 68.50 66.61 51.23 (3.6$\uparrow$)
w/ In-Context Synthesizing Geo3K 26.71 41.24 68.30 67.76 51.00 (3.1$\uparrow$)
w/ Direct Synthesizing Geo3K 26.88 43.53 69.90 68.97 52.32 (5.8$\uparrow$)
w/ Original Questions GeoQA 27.07 43.68 68.90 68.22 51.97 (5.1$\uparrow$)
w/ In-Context Synthesizing GeoQA 26.09 42.87 70.60 69.25 52.20 (5.5$\uparrow$)
w/ Direct Synthesizing GeoQA 26.25 44.64 71.50 68.28 52.67 (6.5$\uparrow$)
w/ Original Questions MMR1 26.15 44.87 72.90 68.74 53.17 (7.5$\uparrow$)
w/ In-Context Synthesizing MMR1 26.15 45.10 71.90 68.62 52.94 (7.0$\uparrow$)
w/ Direct Synthesizing MMR1 26.15 44.11 70.40 67.99 52.16 (5.4$\uparrow$)
@@
To further explore the potential of MM-UPT, we investigate the use of unlabeled synthetic data to improve MLLMs. This aligns with the ultimate goal of MM-UPT: enabling continual self-improvement even after human-created data is exhausted. We utilize two strategies for generating synthetic training samples.
In-Context Synthesizing. Inspired by Self-Instruct [46], we adopt a data generation pipeline based on in-context examples. Each original example includes an image, a question, and its corresponding answer. To synthesize new data, we provide the base MLLM with the full original triplet—image, question, and answer—and ask it to generate a new question that is distinct from the original. During the unsupervised post-training, the MLLM attempts to answer this new question, and we define the pseudo-label as the majority vote among the sampled responses of the model.
Direct Synthesizing. In addition, we also adopt a more straightforward approach to generate synthetic data. Instead of using the full triplet, we prompt the base MLLM with only the image, asking it to create a new question without any context from the original question or answer. This produces a different type of synthetic sample, where the image remains the same, but the question is generated entirely from visual input. As with the in-context approach, the pseudo-label for training is determined by aggregating several model responses through majority voting.
@@
Results. In our experiment, we use the previous two methods to generate the synthetic data, leveraging Geometry3K [23], GeoQA [3], and MMR1 [18] as the seed datasets, and Qwen2.5-VL-7B as the base MLLM for data synthesis. MM-UPT is then applied to the same base model (i.e., Qwen2.5-VL-7B) using each of these different synthetic datasets separately. Table 2 presents experimental results using different synthetic data generation strategies. Both in-context and direct synthesizing lead to significant improvements over the base model, achieving performance comparable to training on original human-written questions. This shows that synthetic questions can effectively enhance the model’s reasoning ability under MM-UPT. Notably, direct synthesizing even surpasses human-written questions (when applied to Geometry3K and GeoQA) on average, demonstrating the strong ability of the model to generate high-quality textual questions solely based on images. This highlights the potential for scalable and fully autonomous self-improvement in multi-modal domain via visual-centric data synthesis.
@@
Further Investigation. Moreover, we manually examine some synthetic data. We observe that in-context synthesizing often produces questions similar to the original ones by substituting conditions or expressions, resembling data rephrasing. In contrast, direct synthesizing generates more diverse and novel questions. While some of the directly synthesized questions still contain hallucinations, many are of high quality and beneficial for unsupervised post-training. This underscores the potential of the direct synthesizing approach as a simple yet effective method for data generation, without the need for textual in-context examples. Below, we present two illustrative examples that showcase the effectiveness and quality of synthetic questions generated through both approaches.
@@
Demo: Examples of synthetic data using different strategies. [Uncaptioned image] Original Question: ${BC}\|{DF}$. $AB=x+5$, $BD=12$, $AC=3x+1$, and $CF=15$. Find $x$. In-Context Synthetic Question: Given that ${BC}\|{DF}$, $AB=2x-3$, $BD=18$, $AC=x+7$, and $CF=24$. Find the value of $x$. Directly Synthetic Question: In the given triangle $\triangle ADF$, point $B$ lies on $AD$ and point $C$ lies on $AF$. If $BC\parallel DF$, what is the ratio of the area of $\triangle ABC$ to the area of $\triangle ADF$?   [Uncaptioned image] Original Question: If $\angle RST$ is a right angle, ${SU}\perp{RT}$, ${UV}\perp{ST},$ and $\angle RTS=47^{\circ},$ find $\angle RSU$ In-Context Synthetic Question: If $\angle RST$ is a right angle, ${SU}\perp{RT}$, ${UV}\perp{ST},$ and $\angle RTS=47^{\circ},$ find $\angle VST$. Directly Synthetic Question: In the given triangle $\triangle RST$, point $U$ lies on $RT$ such that $SU$ is perpendicular to $RT$. Point $V$ lies on $ST$ such that $UV$ is perpendicular to $ST$. If $RU=12$ units, $UT=16$ units, and $SV=9$ units, find the length of $VT$.

### v2

# First SFT, Second RL, Third UPT: Continual Improving Multi-Modal LLM Reasoning via Unsupervised Post-Training

@@ Abstract
[…]
@@ 1 Introduction
Multi-modal Large Language Models (MLLMs) have achieved remarkable performance on a variety of vision-language tasks, ranging from image captioning to visual reasoning [22, 81, 51, 67, 73, 83]. The dominant paradigm for improving MLLMs in the post-training stage typically involves supervised fine-tuning (SFT) and reinforcement learning (RL) [47, 77, 65, 42, 2, 59]. However, both SFT and RL rely on large volumes of high-quality and annotated multi-modal data, such as image captions, visual reasoning traces, verifiable ground truth answers, and human preference signals. As real-world tasks grow in complexity and quantity, a critical challenge emerges: curating and annotating high-quality data at scale becomes increasingly impractical. To overcome this data-dependency, a paradigm shift is required towards a third stage of post-training beyond SFT and RL, dedicated to the continual self-improvement of MLLMs through synthetic and unlabeled data. We formalize this third-stage paradigm as Unsupervised Post-Training.
@@
Previous works have studied the use of MLLMs themselves to generate synthetic instruction data for self-improvement through offline training techniques like SFT and DPO [74, 6, 64, 40, 41, 39, 23]. These approaches typically involve complex pipelines with multiple stages, such as data generation, verification, and filtering, which are hard to iterate online. Fortunately, recent studies demonstrate notable success using online reinforcement learning (e.g. GRPO [37]) with verifiable rewards to enhance the reasoning capabilities of MLLMs [53, 31, 7]. A concurrent work, TTRL [84], further extends this line by applying GRPO on test-time scaling of LLMs. It is promising that online RL enables models to continuously improve, thus acquiring novel reasoning abilities that exceed corresponding base models’ capacity.
@@
Motivated by these insights, we propose MM-UPT (Multi-Modal Unsupervised Post-Training), an easy-to-implement framework for unsupervised post-training in MLLMs. As illustrated in Figure 1, our approach adapts the online reinforcement learning method GRPO [37], which is known for its stability and scalability. The core challenge in applying GRPO to an unsupervised setting is the absence of ground-truth labels for reward calculation. To overcome this, MM-UPT works by deriving implicit reward signals via majority voting over multiple sampled responses. In particular, majority voting aggregates multiple responses and selects the most frequent one, which has been widely used and shown effective to improve model performance [49, 44, 84]. Thus, we adopt the majority-voted answer to serves as a dynamic pseudo-label in GRPO: responses that align with this consensus receive a positive reward, while divergent ones are penalized. This process effectively encourages the model to bootstrap its own high-confidence knowledge, promoting the generation of stable and consistent answers without relying on any external supervision or reward models.
Beyond learning from existing unlabeled data, MM-UPT further extends to a data self-generation setting that enhances scalability under the unsupervised paradigm. Specifically, we design two strategies that allow the MLLM itself to synthesize new training samples: (1) in-context synthesizing, where the model generates new questions conditioned on original examples (image, question, and answer); and (2) direct synthesizing, where the model generates questions solely from the given image. These self-generated samples are then used within the same unsupervised reinforcement learning method, enabling continual self-improvement even in the absence of human-created questions.
@@
[…]
@@
  • We are the first to formalize a three-stage post-training paradigm for MLLMs, with Unsupervised Post-Training (UPT) as the critical third stage enabling continual improvement without external supervision. We instantiate this stage with MM-UPT, a simple and effective framework using majority voting as a self-rewarding signal in online reinforcement learning.
@@
  • We extend MM-UPT to utilize synthetic data generated by the MLLM itself, and find that training the MLLM on such data leads to notable performance gains. This reveals a promising path toward efficient and scalable self-improvement in unsupervised post-training.
@@ 2 Related Works
Self Improvement. High-quality data obtained from human annotations has been shown to significantly boost the performance of LLMs across a wide range of tasks [18, 32, 12]. However, such high-quality annotated data may be exhausted in the future. This presents a substantial obstacle to the continual learning of advanced models. As a result, recent research has shifted toward self-improvement, leveraging data generated by the LLM itself without any external supervision [16, 54, 29, 8, 84]. Several following works also explore self-improvement in the multi-modal domain [11, 6, 64, 74, 39, 72]. Genixer [74] firstly introduces a self-improvement pipeline including complex data generation and filtering for SFT. STIC [6] and SENA [40] construct preference data pairs for DPO [36] in a self-supervised manner, focusing on enhancing perceptual capabilities. In contrast to these approaches which are complex and hard to scale, the key distinction is that our work leverages online reinforcement learning using GRPO [12] with simple yet effective data synthesizing strategies at the post-training stage, which can be more scalable for self-improvement without reliance on any external supervision. In addition, none of these previous methods focus on multi-modal reasoning tasks, which are considered challenging for current models.
Multi-modal Reasoning. Recently, the reasoning abilities of MLLMs have become a central focus of research [27, 82, 53, 56]. In contrast to traditional LLM-based reasoning [28, 62, 12] that primarily relies on text, multi-modal approaches must both process and interpret visual inputs, significantly increasing the complexity of tasks such as geometric problem-solving and chart interpretation [3, 30, 75]. Several works in this field have sought to collect or synthesize a large scale of multi-modal reasoning data [70, 38, 33, 5]. Notably, the recent emergence of o1-like reasoning models [19] represents an initial step toward activating the slow-thinking capabilities of MLLMs, as demonstrated by several SFT-based methods, such as LLaVA-CoT [55], MAmmoTH-VL [13], and Mulberry [60]. Moreover, some concurrent works have further explored reinforcement learning approaches, particularly GRPO [37], in the post-training stage of MLLMs to enhance performance on multi-modal reasoning tasks [78, 31, 7, 48, 34]. While these supervised post-training methods have demonstrated promising results, our work explores a different direction by focusing on totally unsupervised post-training of MLLMs to self-improve the reasoning abilities.
@@ 3 The Framework of Multi-Modal Unsupervised Post-Training
Existing post-training techniques for MLLMs, such as supervised fine-tuning (SFT) and reinforcement learning (RLHF or RLVR), rely heavily on labeled data or external reward models. While these approaches have proven effective, their reliance on external supervision makes continual improvement unsustainable due to the high cost and limited scalability of manual annotation. To overcome this limitation, we formalize a new third-stage post-training paradigm that enables the model to self-improve without access to any external supervision, such as ground-truth labels or additional reward models. We instantiate this paradigm with MM-UPT (Multi-Modal Unsupervised Post-Training), a simple yet effective framework designed to operate purely on unlabeled multi-modal data. The overview of our complete framework is shown in Figure 1.
@@ 3.1 Problem Formulation
3.2 Training Method
To achieve the unsupervised training, MM-UPT introduces a self-rewarding mechanism using majority voting as pseudo-labels [49] based on the online reinforcement learning. In particular, MM-UPT is built upon the GRPO algorithm [37], which is widely used in the post-training stage of multi-modal LLMs. GRPO optimizes computational efficiency by eliminating the need for a separate value model; instead, it directly utilizes group-normalized rewards to estimate advantages. Specifically, for a question $q$ and the correlated image $I$ from the training dataset $Q$, GRPO samples a group of responses $O=\{o_{i}\}_{i=1}^{G}$ from the old policy $\pi_{old}$ and then optimizes the policy model by maximizing the following objective:
@@ 3.2 Training Method
Algorithm 1 The training method of MM-UPT
@@
In this way, we compute pseudo-rewards via majority voting and apply standard GRPO to update the MLLM. This majority-based reward encourages the model to converge toward consistent, high-consensus responses, thereby enabling the model to further exploit its existing self-knowledge leveraging unlabeled data. The pipeline of our training method in MM-UPT is shown in Algorithm 1.
3.3 Synthetic Data
To further explore the scalability of unsupervised post-training, we extend our framework to a data self-generation setting, where the model is asked to synthesize new training samples on its own. In particular, we design two simple yet effective data synthesizing strategies as follows.
In-Context Synthesizing. Inspired by Self-Instruct [50], we construct a data generation pipeline that leverages in-context examples to guide the synthesis process. Each original example consists of an image, a question, and its corresponding answer. To synthesize new samples, we provide the model with the full triplet and instruct it to generate a new question that is semantically distinct from the original but relevant to the same image. This strategy helps generate task-relevant and meaningful variations of the original question, as well as ensure the quality of synthetic questions. During unsupervised post-training, the model then attempts to answer each newly generated question, and pseudo-labels are derived from the majority vote among its sampled responses, consistent with the mechanism described in Section 3.
Direct Synthesizing. In addition to in-context generation, we also explore a direct synthesizing strategy that further increases diversity. Here, the model receives only the image and is prompted to freely create a new question without any reference to the original question. This open-ended formulation encourages the model to generate a wider range of diverse and novel questions based solely on the visual input, rather than being constrained by the original task. Similar to the in-context setup, we perform unsupervised post-training on these synthetic samples using majority voting to define pseudo-rewards.
This setup enables the MLLM to expand the training corpus without any human annotations, thus achieving a fully autonomous self-improvement loop.
@@ 4 Experiments
Table 1: Main results of Scenario 1 on four multi-modal mathematical reasoning benchmarks. We report accuracy (%) for each method on MathVision, MathVerse, MathVista, and We-Math. All methods are conducted on the Qwen2.5-VL-7B backbone. $M$-UPT outperforms other baseline methods, and is even competitive with supervised methods.
@@
+ MM-UPT (Ours) Geometry3K 27.33 42.46 68.50 66.61 51.23
+ MM-UPT (Ours) GeoQA 27.07 43.68 68.90 68.22 51.97
+ MM-UPT (Ours) MMR1 26.15 44.87 72.90 68.74 53.17
@@ 4.1 Experimental Setup
Baseline Methods. Several prior works have explored self-improvement in both LLMs and MLLMs. Note that we focus on unsupervised self-improvement, we do not compare with methods that rely on external models (e.g., GPT-4o [18]) for supervision [71, 17, 79, 80, 64]. Instead, we compare with several totally unsupervised methods: LMSI [16], SRLM [54], Genixer [74], and STIC [6]. In particular, LMSI corresponds to supervised fine-tuning with self-generated content selected by majority voting. SRLM uses the model itself as LLM-as-a-Judge [76] to provide its own rewards during DPO [36] training. Genixer prompts the MLLM to first self-generate an answer and then self-check it. STIC applies DPO where original images and good prompts are used to generate preferred answers, and corrupted images and bad prompts to produce rejected answers. Additionally, we also compare with GRPO [37] and rejection SFT [43], which are two strong supervised methods. The details of these baseline methods are shown in Appendix A.1.
@@
Implementation Details. We adopt the EasyR1 [61] framework for multi-modal unsupervised post-training, which is based on GRPO. Specifically, we set the training episodes to 15, and use AdamW optimizer [24] with a learning rate of $1\times 10^{-6}$, weight decay of $1\times 10^{-2}$, and gradient clipping at a maximum norm of 1.0. The KL divergence constraint $\beta$ in GRPO is set to 0.01 to stabilize the training. The vision tower of the multi-modal model is also tuned without freezing. In our training, we use a rollout temperature of 0.7, which strikes a good balance: lower temperatures produce low-diversity outputs, while higher temperatures often lead to lower-quality outputs. Within each episode, we perform one rollout group (GG=10) per data point. Other hyperparameters follow the default settings provided in the EasyR1 framework.
4.2 Scenario 1: Unsupervised Training on Standard Datasets
@@
Table 1 presents the main results on four challenging multi-modal mathematical reasoning benchmarks. We observe that MM-UPT achieves consistent improvements in average over the base Qwen2.5-VL-7B model across all datasets, also outperforming other baseline methods such as SRLM, LMSI, Genixer, and STIC. Notably, MM-UPT is able to improve the average score from 49.47 (base model) to 53.17 (with MMR1 dataset), demonstrating its effectiveness in leveraging unlabeled data for self-improvement. In comparison, previous baselines provide only marginal gains or even degrade performance on certain benchmarks, highlighting the limitations of existing methods when applied to already strong models in multi-modal reasoning tasks. Furthermore, we find that MM-UPT is even competitive with supervised post-training methods, such as rejection sampling-based SFT [43] and GRPO [37]. These results underscore the potential of MM-UPT to further exploit the knowledge embedded in multi-modal models for self-improvement. Further experiments on the training dynamics, generalization and adaptability of our method are shown in Appendix B.
4.3 Scenario 2: Unsupervised Training on Synthetic Datasets
Table 2: Performance comparison of MM-UPT using different synthetic data generation strategies in Scenario 2. Both “In-Context Synthesizing” and “Direct Synthesizing” approaches yield significant improvements over the base model and perform competitively with the “Original Questions” on average, demonstrating the effectiveness of synthetic data for unsupervised self-improvement.
@@ 4.3 Scenario 2: Unsupervised Training on Synthetic Datasets
[…]
@@
To further explore the potential of MM-UPT, we investigate the use of unlabeled synthetic data (mentioned in Section 3.3) to improve MLLMs. This aligns with the ultimate goal of MM-UPT: enabling continual self-improvement even after human-created data is exhausted.
@@
In our experiment, we use the previous two methods in Section 3.3 to generate the synthetic data, leveraging Geometry3K [25], GeoQA [4], and MMR1 [20] as the seed datasets, and Qwen2.5-VL-7B as the base MLLM for data synthesis. MM-UPT is then applied to the same base model (i.e., Qwen2.5-VL-7B) using each of these different synthetic datasets separately. Table 2 presents experimental results using different synthetic data generation strategies. Both in-context and direct synthesizing lead to significant improvements over the base model, achieving performance comparable to training on original human-written questions. This shows that synthetic questions can effectively enhance the model’s reasoning ability under MM-UPT. Notably, direct synthesizing even surpasses human-written questions (when applied to Geometry3K and GeoQA) on average, demonstrating the strong ability of the model to generate high-quality textual questions solely based on images. This highlights the potential for scalable and fully autonomous self-improvement in multi-modal domain via visual-centric data synthesis.
@@
Moreover, we manually examine some synthetic data. We observe that in-context synthesizing often produces questions similar to the original ones by substituting conditions or expressions, resembling data rephrasing. In contrast, direct synthesizing generates more diverse and novel questions. While some of the directly synthesized questions still contain hallucinations, many are of high quality and beneficial for unsupervised post-training. This underscores the potential of the direct synthesizing approach as a simple yet effective method for data generation, without the need for textual in-context examples. Below, we present two illustrative examples that showcase the effectiveness and quality of synthetic questions generated through both approaches.
@@
Demo: Examples of synthetic data using different strategies. [Uncaptioned image] Original Question: ${BC}\|{DF}$. $AB=x+5$, $BD=12$, $AC=3x+1$, and $CF=15$. Find $x$. In-Context Synthetic Question: Given that ${BC}\|{DF}$, AB=2x3AB=2x-3, $BD=18$, $AC=x+7$, and $CF=24$. Find the value of $x$. Directly Synthetic Question: In the given triangle $\triangle ADF$, point $B$ lies on ADAD and point $C$ lies on AFAF. If $BC\parallel DF$, what is the ratio of the area of $\triangle ABC$ to the area of $\triangle ADF$?   [Uncaptioned image] Original Question: If $\angle RST$ is a right angle, ${SU}\perp{RT}$, ${UV}\perp{ST},$ and $\angle RTS=47^{\circ},$ find $\angle RSU$ In-Context Synthetic Question: If $\angle RST$ is a right angle, ${SU}\perp{RT}$, ${UV}\perp{ST},$ and $\angle RTS=47^{\circ},$ find $\angle VST$. Directly Synthetic Question: In the given triangle $\triangle RST$, point $U$ lies on $RT$ such that $SU$ is perpendicular to $RT$. Point $V$ lies on $ST$ such that $UV$ is perpendicular to $ST$. If $RU=12$ units, $UT=16$ units, and $SV=9$ units, find the length of $VT$.